# To Label or Not to Label: Hybrid Active Learning for Neural Machine Translation

### v1

@@ Abstract
Active learning (AL) techniques reduce labeling costs for training neural machine translation (NMT) models by selecting smaller representative subsets from unlabeled data for annotation. Diversity sampling techniques select heterogeneous instances, while uncertainty sampling methods select instances with the highest model uncertainty. Both approaches have limitations - diversity methods may extract varied but trivial examples, while uncertainty sampling can yield repetitive, uninformative instances. To bridge this gap, we propose HUDS, a hybrid AL strategy for domain adaptation in NMT that combines uncertainty and diversity for sentence selection. HUDS computes uncertainty scores for unlabeled sentences and subsequently stratifies them. It then clusters sentence embeddings within each stratum using k-MEANS and computes diversity scores by distance to the centroid. A weighted hybrid score that combines uncertainty and diversity is then used to select the top instances for annotation in each AL iteration. Experiments on multi-domain German-English datasets demonstrate the better performance of HUDS over other strong AL baselines. We analyze the sentence selection with HUDS and show that it prioritizes diverse instances having high model uncertainty for annotation in early AL iterations.
@@
Abdul Hameed Azeemi and Ihsan Ayyub Qazi and Agha Ali Raza Lahore University of Management Sciences {abdul.azeemi, ihsan.qazi, agha.ali.raza}@lums.edu.pk
@@ 1 Introduction
The training process for Machine Translation (MT) models typically requires large, high-quality parallel corpora to achieve robust performance. However, acquiring such data is costly and laborious, especially for low-resource languages and domain-specific data, requiring professional translators with domain knowledge. For instance, training an MT model to translate specific medical conditions requires extensive labeling by medicine experts. Active learning (AL) is a well-known strategy that helps to reduce the labeling requirements by selecting a smaller representative subset for annotation, thereby reducing its overall cost. AL follows an iterative procedure that involves (1) querying instances from an unlabeled pool; (2) obtaining annotations for unlabeled instances through a human expert; and (3) adding the annotated examples to the labeled dataset and retraining.
@@
In the context of Neural Machine Translation (NMT), active learning aids in selecting a small number of sentences for labeling that are likely to bring a similar performance if the NMT model is trained on a larger labeled set. AL methods for NMT can be broadly classified into diversity and uncertainty sampling methods, often described as the "two faces of active learning" Dasgupta (2011). Diversity sampling ensures that a heterogeneous set of instances from the unlabeled set is selected for annotation. In contrast, uncertainty sampling attempts to select the instances with the highest model uncertainty, i.e., the ones which are the most difficult for the next AL iteration. Both these methods have their limitations, such that the methods focusing on diversity can extract varied but trivial examples, whereas uncertainty sampling may lead to the selection of instances with high uncertainty but are repetitive and not useful. To mitigate this, hybrid sampling approaches have been proposed for other NLP tasks, including named entity recognition Kim et al. (2006), abstractive text summarization Tsvigun et al. (2022a), and text classification Yuan et al. (2020). For NMT, Hu and Neubig (2021) propose an AL strategy for NMT to select informative sentences and phrases for annotation, but the method does not consider model uncertainty scores while sampling. No hybrid AL strategy for efficiently acquiring domain-specific data in NMT has been proposed yet which successfully incorporates model uncertainty and data diversity in the sampling procedure.
@@
Figure 1: A conceptual representation of Hybrid Uncertainty and Diversity Sampling (HUDS) for AL in NMT.
@@
To develop a hybrid AL strategy for NMT, we should consider a multi-step procedure involving uncertainty computation, diversity sampling, and a weighted hybrid acquisition function that leverages both. In this work, we develop HUDS (Hybrid Uncertainty and Diversity Sampling), a hybrid active learning strategy for NMT. We first compute uncertainty scores for unlabeled sentences and stratify them. We then obtain embeddings for each sentence in a stratum, cluster them using $k$-MEANS and compute the diversity score for each instance by their distance to the centroid. Finally, we compute a hybrid sampling score through a weighted sum of the uncertainty and diversity scores and select the top-k instances with the highest scores for annotation in each iteration of AL. This weighted sampling ensures the selection of diverse yet challenging instances for annotation, which contributes to the model’s robust performance on domain-specific datasets. We evaluate our approach on three corpora spanning different domains (medicine, law, IT) with the BART language model and demonstrate that HUDS outperforms other AL baselines in SacreBLEU score. Finally, we examine the examples selected for annotation by HUDS and demonstrate their appropriateness for AL.
@@ 2.1 Neural Machine Translation
$\mathcal{L}_{\mathrm{CE}}(\theta)=-\sum_{i=1}^{N}\log p\left(y_{i}\mid y_{0:i},\mathbf{X};\theta\right)$ (2)
@@ 2.2 Active Learning for NMT
We consider a labeled corpus $(x,y)\in\mathcal{D}_{l}$ and an unlabeled corpus $(x,y)\in\mathcal{D}_{u}$ for active learning in NMT. In the context of domain adaptation, $\mathcal{D}_{u}$ is a larger out-of-domain labeled corpus (e.g., WMT14 German-English corpus), and $\mathcal{D}_{l}$ is a small in-domain labeled corpus (e.g., Medicine domain German-English corpus). For acquiring the translation of the sentences selected by the query strategy, we consider a human oracle $\mathcal{O}(\cdot)$ who can translate any sentence in the unlabeled data $\mathcal{O}(x)\forall x\in\mathcal{D}_{u}$. We consider a fixed budget $c$ for annotation such that all the annotation through $\mathcal{O}$ must occur within that budget.
Input: Unlabeled pool $\mathcal{U}$, Number of strata $n$, Pre-trained encoder $\mathcal{P}$, Initial Model $\mathcal{M}$, Hybrid parameter $\lambda$
1 $S\leftarrow\mathcal{M}(u)$, $u\in\mathcal{U}$triangleright Compute uncertainty scores for each sentence in $\mathcal{U}$ $s_{min}\leftarrow min(S)$ $s_{max}\leftarrow max(S)$ $r\leftarrow s_{max}-s_{min}$ $\mathcal{N}\leftarrow\emptyset$ for $i=1\to n$ do
2       $\mathcal{N}\leftarrow\mathcal{N}\cup{\mathcal{S}}\left(s_{\text{min}}+\frac{i-1}{n}(r),s_{\text{min}}+\frac{i}{n}(r)\right)$triangleright Perform stratification on uncertainty scores ${\mathcal{S}}$
3 end for
[…]
5       $\mathcal{E}\leftarrow\emptyset$ for sentences $x\in\mathcal{X}$ do
6             $\mathcal{E}\leftarrow\mathcal{E}\cup\mathcal{P}(x)$triangleright Obtain embeddings for each sentence in stratum $\mathcal{X}$
7       end for
[…]
9 end for
$\mathcal{Q}\leftarrow topK(\mathcal{C})$triangleright Select top k instances for annotation Output: $\mathcal{Q}$
Algorithm 1 Single iteration of HUDS
@@ 2.3.1 Badge
The primary objective of the BADGE algorithm is to select heterogeneous and uncertain instances for annotation in each iteration of AL. This is achieved by clustering special representations called gradient embeddings that encapsulate model certainty. It is the gradient of the cross-entropy loss with respect to the model’s last layer parameters. For an unlabeled instance $x$, the $i$-th block of gradient embedding is represented as:
$\left(g_{x}\right)_{i}=\frac{\partial}{\partial W_{i}}\ell_{\mathrm{CE}}(f(x;\theta),\hat{y})$ (3)
where $\hat{y}$ is the label predicted for $x$ by the trained model.
Once the gradient embeddings are computed, BADGE uses $k$-MEANS++ seeding algorithm Arthur and Vassilvitskii (2007) to select a subset of examples $S_{t}$ for annotation. The examples included in this subset are diverse in terms of uncertainty and data distribution due to the way gradient embeddings are calculated.
@@ 2.3.2 ALPS
Improving upon Badge, Alps attempts to estimate the uncertainty of instances using the masked language model (MLM) loss of a pre-trained BERT within the downstream classification task. Thus, Alps works in a cold-start setting by leveraging the loss from a pre-trained BERT in contrast to Badge which estimates uncertainty using the classification model that is being trained, i.e, in a warm-start setting. Alps achieves this by creating a surprisal embedding $s_{x}$ for each input $x$. This process involves feeding the non-masked input $x$ into the BERT MLM head and calculating the cross-entropy loss for 15% randomly selected tokens. These embeddings are then clustered and a subset of examples are selected for the next iteration.
@@ 3 Hybrid Uncertainty and Diversity Sampling
3.1 Uncertainty scores
We first estimate the uncertainty for an unlabeled sentence $s$ through the normalized negative log-likelihood (NNLL).
$NNLL(s)=-\frac{1}{T}\sum_{t=1}^{T}\log p(w_{t}|w_{1},w_{2},\dots,w_{t-1})$ (4)
where $T$ is the sentence length and $w_{t}$ is the $t^{th}$ token in $s$. Lower NLL values indicate higher model confidence in the prediction, while higher values indicate less confidence thus greater uncertainty. We then perform stratification on uncertainty scores, with the range for $i$th stratum defined as,
@@ 3.1 Uncertainty scores
3.2 Diversity scores
We then utilize a pre-trained BERT to obtain the embeddings of sentences within each stratum. Subsequently, we cluster the embeddings using $k$-MEANS and compute the diversity score for each instance by the cosine distance between that instance and the cluster centroid. The goal of clustering is only to determine the distance of each example from the centroid of the (already defined) stratum and thus a value of $k=1$ suffices here. These diversity scores represent the heterogeneity of the examples. Therefore, the semantically dissimilar (and potentially harder) instances in each stratum have higher scores than the prototypical examples.
3.3 Hybrid scores
We compute a hybrid sampling score $H(x)$ for a sentence $x$ through the weighted sum of the uncertainty and diversity scores (Eq. 6) and select the top-k instances with the highest scores for annotation.
@@ 3.3 Hybrid scores
Stratification of uncertainty scores allows the selection of diverse sentences from each subpopulation, i.e., examples with a low uncertainty can also be selected as long as they are diverse. This is in contrast to the other pure uncertainty sampling methods which prevent the selection of informative sentences that have a low uncertainty. We hypothesize that this aids in selecting sentences with low overall uncertainty but having harder (dissimilar) segments that the NMT model cannot translate well.
@@ 4 Experimental Setup
We mirror the conventional active learning setup followed in previous studies Shelmanov et al. (2021); Dor et al. (2020); Hu and Neubig (2021); Tsvigun et al. (2022a). We first train the NMT model on the out-of-domain labeled corpus. We then start the AL procedure with a randomly selected small subset of labeled sentences from in-domain data and in the first iteration fine-tune the NMT model on this subset. In each subsequent iteration, we use the query strategy $\mathcal{Q}$ to select the examples for annotation, add them to the annotated pool of examples, and remove them from the unlabeled set. The NMT model is then trained on the sentence pairs in the labeled pool and evaluated on an unseen validation set. We use SacreBLEU Post (2018) to evaluate the performance on the validation set in each iteration.
@@ 4.1 Datasets
We utilize the WMT14 En-De (English-German) parallel data as our out-of-domain labeled dataset. It contains 4.5M training pairs and 3000 validation and test pairs.
For the in-domain data, we use the multi-domains dataset containing German-English parallel data for five domains Koehn and Knowles (2017), deduplicated and resplit by Aharoni and Goldberg (2020). This dataset has been frequently used for the evaluation of domain adaptation and active learning approaches in neural machine translation Dou et al. (2019); Hu and Neubig (2021) We use three corpora for our experiments: Medicine, Law, and IT. The statistics are shown in Table 1.
@@ 4.2 Model
We use the BART model [^1] https://huggingface.co/facebook/bart-base for our experiments which works well for natural language generation tasks including translation Lewis et al. (2020). BART is a transformer encoder-decoder model having a bidirectional encoder and an autoregressive decoder. We conduct experiments using the BART-base variant pre-trained on Wikipedia and the BookCorpus. This variant has 140 million parameters and six blocks in the encoder and decoder. We train the model using the AdamW optimizer with a training batch size of 16. The learning rate is set to 2e-5 with a weight decay of 0.028 and gradient clipping of 0.28. The beam size is set to 4 for evaluation. We use $\lambda=0.5$ for HUDS sampling, giving equal weightage to uncertainty and diversity in the computation of hybrid scores. Uncertainty is computed for 20,000 randomly selected examples from the unlabeled pool as it is computationally expensive to be performed for the complete dataset. The experiments are run on a single 48GB NVIDIA A6000 GPU.
(a) Medicine Dataset
(b) Law Dataset
(c) IT Dataset
Figure 2: BLEU scores with different active learning strategies for NMT across different domain datasets. BART-base model fine-tuned with WMT14 En-De dataset is used in all experiments. Active learning is done over ten iterations with 1000 sentences queried in each iteration. For each strategy, three independent runs are done with different seeds, and the mean SacreBLEU is reported, with the shaded regions representing the standard deviation. HUDS consistently shows better performance compared to other AL strategies.

### v2

@@ Abstract
Active learning (AL) techniques reduce labeling costs for training neural machine translation (NMT) models by selecting smaller representative subsets from unlabeled data for annotation. Diversity sampling techniques select heterogeneous instances, while uncertainty sampling methods select instances with the highest model uncertainty. Both approaches have limitations - diversity methods may extract varied but trivial examples, while uncertainty sampling can yield repetitive, uninformative instances. To bridge this gap, we propose Hybrid Uncertainty and Diversity Sampling (HUDS), an AL strategy for domain adaptation in NMT that combines uncertainty and diversity for sentence selection. HUDS computes uncertainty scores for unlabeled sentences and subsequently stratifies them. It then clusters sentence embeddings within each stratum and computes diversity scores by distance to the centroid. A weighted hybrid score that combines uncertainty and diversity is then used to select the top instances for annotation in each AL iteration. Experiments on multi-domain German-English and French-English datasets demonstrate the better performance of HUDS over other strong AL baselines. We analyze the sentence selection with HUDS and show that it prioritizes diverse instances having high model uncertainty for annotation in early AL iterations.
@@
Abdul Hameed Azeemi  and Ihsan Ayyub Qazi  and Agha Ali Raza Lahore University of Management Sciences Pakistan {abdul.azeemi, ihsan.qazi, agha.ali.raza}@lums.edu.pk
@@ 1 Introduction
Figure 1: Comparison of the instances selected through different active learning strategies on the task of English to German (En $\rightarrow$ De) translation using the Law domain dataset and BART-base model. The color represents the active learning iteration; from the darkest (first iteration) to the lightest (tenth iteration). Our proposed hybrid active learning method (HUDS) selects diverse yet challenging sentences for annotation in early iterations.
@@
The training process for Machine Translation (MT) models typically requires large, high-quality parallel corpora to achieve robust performance. However, acquiring such data is costly and laborious, especially for low-resource languages and domain-specific data which requires professional translators with domain knowledge. For instance, training an MT model to translate specific medical conditions requires extensive labeling by medicine experts. Active learning (AL) is a well-known strategy that helps to reduce the labeling requirements by selecting a smaller representative subset for annotation, thereby reducing the overall cost. AL follows an iterative procedure that involves (1) querying instances from an unlabeled pool; (2) obtaining annotations for unlabeled instances through a human expert; and (3) adding the annotated examples to the labeled dataset and retraining.
@@
In the context of Neural Machine Translation (NMT), active learning aids in selecting a small number of sentences for labeling that are likely to bring a similar performance if the NMT model is trained on a larger labeled set. AL methods for NMT can be broadly classified into diversity and uncertainty sampling methods, often described as the “two faces of active learning” Dasgupta (2011). Diversity sampling ensures that a heterogeneous set of instances from the unlabeled set is selected for annotation. In contrast, uncertainty sampling attempts to select the instances with the highest model uncertainty, i.e., the ones which are the most difficult for the next AL iteration. Both these methods have their limitations, such that the methods focusing on diversity can extract varied but trivial examples, whereas uncertainty sampling may lead to the selection of instances with high uncertainty but which are repetitive and not useful. To mitigate this, hybrid sampling approaches have been proposed for other NLP tasks, including named entity recognition Kim et al. (2006), abstractive text summarization Tsvigun et al. (2022a), and text classification Yuan et al. (2020). No hybrid AL strategy for efficiently acquiring domain-specific data in NMT has been proposed yet which successfully incorporates model uncertainty and data diversity in the sampling procedure.
@@
Figure 2: A representation of hybrid uncertainty and diversity sampling (HUDS) for active learning in NMT.
@@
To develop a hybrid AL strategy for NMT, we should consider a multi-step procedure involving uncertainty computation, diversity sampling, and a weighted hybrid acquisition function that leverages both. In this work, we develop HUDS (Hybrid Uncertainty and Diversity Sampling), a hybrid active learning strategy for NMT. We first compute uncertainty scores for unlabeled sentences and stratify them. We then obtain embeddings for each sentence in a stratum, cluster them using $k$-MEANS and compute the diversity score for each instance by their distance to the centroid. Finally, we compute a hybrid sampling score through a weighted sum of the uncertainty and diversity scores and select the top-k instances with the highest scores for annotation in each iteration of AL. This weighted sampling ensures the selection of diverse yet challenging instances for annotation (Fig. 1), which contributes to the model’s robust performance on domain-specific datasets.
1.1 Contributions
  • To the best of our knowledge, we present the first hybrid active learning strategy (HUDS) for NMT.
  • We evaluate HUDS on German-English and French-English datasets spanning different domains (Medicine, Law, IT, Ted talks) with the BART language model and demonstrate that HUDS outperforms other AL baselines in NMT.
  • We examine the examples selected for annotation by HUDS and demonstrate that it prioritizes diverse yet challenging instances for annotation in early AL iterations. We also find that HUDS selects varied instances with higher unigram coverage compared to other AL strategies.
@@ 2.1 Neural Machine Translation
$\mathcal{L}_{\mathrm{CE}}(\theta)=-\sum_{i=1}^{T}\log p\left(y_{i}\mid y_{0:i},\mathbf{X};\theta\right)$ (2)
@@ 2.2 Active Learning for NMT
We consider a labeled corpus $\mathcal{D}_{l}$ and an unlabeled corpus $\mathcal{D}_{u}$ for active learning in NMT. In the context of domain adaptation, $\mathcal{D}_{l}$ is a larger out-of-domain labeled corpus (e.g., WMT14 German-English corpus), and $\mathcal{D}_{u}$ is a small in-domain unlabeled corpus (e.g., Medicine domain German-English corpus). For acquiring the translation of the sentences selected by the query strategy, we consider a human oracle who can translate any sentence in the unlabeled data. We consider a fixed budget $c$ for annotation such that all the annotation through must occur within that budget.
@@ 2.3.1 Badge
The primary objective of the BADGE algorithm is to select heterogeneous and uncertain instances for annotation in each iteration of AL. This is achieved by clustering special representations called gradient embeddings that encapsulate model certainty. Once the gradient embeddings are computed, BADGE uses $k$-MEANS++ seeding algorithm Arthur and Vassilvitskii (2007) to select a subset of examples $S_{t}$ for annotation. The examples included in this subset are diverse in terms of uncertainty and data distribution due to the way gradient embeddings are calculated.
@@ 2.3.2 ALPS
Improving upon Badge, Alps attempts to estimate the uncertainty of instances using the masked language model (MLM) loss of a pre-trained BERT within the downstream classification task. Thus, Alps works in a cold-start setting by leveraging the loss from a pre-trained BERT in contrast to Badge which estimates uncertainty using the classification model that is being trained, i.e, in a warm-start setting. Alps achieves this by creating a surprisal embedding $s_{x}$ for each input $x$. This process involves feeding the non-masked input $x$ into the BERT MLM head and calculating the cross-entropy loss for 15% randomly selected tokens. These embeddings are then clustered and a diverse subset of examples are selected for the next iteration.
Input: Unlabeled pool $\mathcal{D}_{u}$, Number of strata $n$, Pre-trained encoder $\mathcal{P}$, Acquisition Model $\mathcal{M}$, Hybrid parameter $\lambda$, Instances to select $k$
$\mathcal{U}_{s}\leftarrow\mathcal{M}(u)$, $u\in\mathcal{D}_{u}$triangleright Compute uncertainty scores for each sentence in $\mathcal{D}_{u}$
$s_{min}\leftarrow min(\mathcal{U}_{s})$, $s_{max}\leftarrow max(\mathcal{U}_{s})$, $r\leftarrow s_{max}-s_{min}$, $\mathcal{N}\leftarrow\emptyset$
for $i=1\to n$ do
       $\mathcal{N}\leftarrow\mathcal{N}\cup\mathcal{U}_{s}\left(s_{\text{min}}+\frac{i-1}{n}(r),s_{\text{min}}+\frac{i}{n}(r)\right)$triangleright Perform stratification on uncertainty scores $\mathcal{U}_{s}$
end for
$\mathcal{C}\leftarrow\emptyset$
for $\mathcal{X}$ in $\mathcal{N}$ do
      $\mathcal{E}\leftarrow\emptyset$
      for sentences $x\in\mathcal{X}$ do
             $\mathcal{E}\leftarrow\mathcal{E}\cup\mathcal{P}(x)$triangleright Obtain embeddings for each sentence in stratum $\mathcal{X}$
       end for
[…]
       $\triangleright$ Perform $k$-MEANS clustering on embeddings
      $\mathcal{C}\leftarrow\mathcal{C}\cup\lambda\cdot d(c,e)+(1-\lambda)\cdot\mathcal{U}_{s}(e),e\in\mathcal{E}$
       $\triangleright$ Compute hybrid score for each sentence
end for
$\mathcal{Q}\leftarrow topK(\mathcal{C},k)$triangleright Select top k instances for annotation
Output: $\mathcal{Q}$
Algorithm 1 Single iteration of HUDS
@@ 3 Hybrid Uncertainty and Diversity Sampling
Uncertainty scores. We first estimate the uncertainty for an unlabeled sentence $s$ through the normalized negative log-likelihood: $NNLL(X)=-\frac{1}{S}\sum_{j=1}^{S}\log p(x_{j}|x_{1},x_{2},\dots,x_{j-1})$, where $S$ is the sentence length and $x_{j}$ is the $j^{th}$ token in $X$. Lower NLL values indicate higher model confidence in the prediction, while higher values indicate less confidence thus greater uncertainty. We then perform stratification on uncertainty scores, with the range for $i$th stratum defined as,
@@
Diversity scores. We then utilize a pre-trained BERT to obtain the embeddings of sentences within each stratum. Subsequently, we cluster the embeddings using $k$-MEANS and compute the diversity score for each instance by the cosine distance between that instance and the cluster centroid. The goal of clustering is only to determine the distance of each example from the centroid of the (already defined) stratum and thus a value of $k=1$ suffices here. In other words, diversity is represented with distance from the average vector, over sentences grouped by uncertainty strata. These diversity scores represent the heterogeneity of the examples. Therefore, the semantically dissimilar (and potentially harder) instances in each stratum have higher scores than the prototypical examples.
Hybrid scores. We compute a hybrid sampling score $H(x)$ for a sentence $x$ through the weighted sum of the uncertainty and diversity scores (Eq. 4) and select the top-k instances with the highest scores for annotation.
@@
Stratification of uncertainty scores allows the selection of diverse sentences from each subpopulation, i.e., examples with a low uncertainty can also be selected as long as they are diverse. This is in contrast to the other pure uncertainty sampling methods which prevent the selection of informative sentences that have a low uncertainty. We hypothesize that hybrid sampling aids in selecting sentences with low overall uncertainty but having harder (dissimilar) segments that the NMT model cannot translate well.
@@ 4 Experimental Setup
We mirror the conventional active learning setup followed in previous studies Shelmanov et al. (2021); Dor et al. (2020); Hu and Neubig (2021a); Tsvigun et al. (2022a). We first train the NMT model on the out-of-domain labeled corpus. We then start the AL procedure with a randomly selected small subset of labeled sentences from in-domain data and in the first iteration fine-tune the NMT model on this subset. In each subsequent iteration, we use the query strategy $\mathcal{Q}$ to select the examples for annotation, add them to the annotated pool of examples (i.e., emulate the manual labeling process of translation without human involvement), and remove them from the unlabeled set. The NMT model is then trained on the sentence pairs in the labeled pool and evaluated on an unseen validation set. We use SacreBLEU Post (2018) to evaluate the performance on the validation set in each iteration, according to the protocol followed in previous works on AL in NMT Hu and Neubig (2021b); Zhao et al. (2020). SacreBLEU uses the standard WMT tokenization, addresses several issues in the reporting of regular BLEU scores, and is more reproducible Post (2018).
@@ 4.1 Datasets
We utilize the WMT14 En-De (English-German) parallel data as our out-of-domain labeled dataset. It contains 4.5M training pairs and 3000 validation and test pairs. For the in-domain data, we use the multi-domains dataset containing German-English parallel data for five domains Koehn and Knowles (2017), deduplicated and resplit by Aharoni and Goldberg (2020). This dataset has been frequently used for the evaluation of domain adaptation and active learning approaches in neural machine translation Dou et al. (2019); Hu and Neubig (2021a). We use three corpora for our experiments: Medicine, Law, and IT. The statistics are shown in Table 1.
@@ 4.2 Model
We use the BART model [^1] https://huggingface.co/facebook/bart-base for our experiments which works well for natural language generation tasks including translation Lewis et al. (2020). BART is a transformer encoder-decoder model having a bidirectional encoder and an autoregressive decoder. Our choice of BART is motivated by recent works on active learning for text generation that have employed encoder-decoder architectures such as BART and demonstrated their effectiveness Xia et al. (2024); Tsvigun et al. (2022a). We did not use a large language model (LLM) for the domain adaptation experiments, since LLMs are typically pre-trained on massive, diverse corpora that often overlap with the test domain. Such pre-training raises the risk of data contamination Sainz et al. (2023), making it difficult to accurately assess the model’s ability to learn from newly introduced, domain-specific data. As a result, improvements observed in an LLM would not reliably reflect genuine gains from the active learning process, but rather the model’s prior exposure to similar samples.
We conduct experiments using the BART-base variant pre-trained on Wikipedia and the BookCorpus. This variant has 140 million parameters and six blocks in the encoder and decoder. We train the model using the AdamW optimizer with a training batch size of 16. The learning rate is set to 2e-5 with a weight decay of 0.028 and gradient clipping of 0.28. The beam size is set to 4 for evaluation. We use $\lambda=0.5$ for HUDS sampling (after the first query iteration), giving equal weightage to uncertainty and diversity in the computation of hybrid scores. Uncertainty is computed for 20,000 randomly selected examples from the unlabeled pool as it is computationally expensive to be performed for the complete dataset. 1000 sentences are queried in each iteration. The experiments are run on a single 48GB NVIDIA A6000 GPU.
Dataset AL AL Iteration
Strategy 2 4 6 8 10
Medicine Random 29.36±0.1 30.75±0.1 31.44±0.1 31.93±0.1 32.70±0.1
NSP 29.69±0.6 31.39±0.5 32.27±0.4 32.89±0.3 33.68±0.2
IDDS 29.45±0.1 31.11±0.1 32.08±0.3 32.76±0.2 33.05±0.0
HUDS 30.76±0.2 32.03±0.1 33.14±0.1 33.65±0.1 34.58±0.1
IT Random 27.64±0.3 28.30±0.1 28.78±0.1 29.28±0.1 29.54±0.1
NSP 28.43±0.2 29.26±0.1 29.70±0.2 30.09±0.1 30.23±0.2
IDDS 23.81±0.1 24.31±0.2 24.84±0.0 25.21±0.1 25.41±0.1
HUDS 29.15±0.1 29.69±0.1 30.12±0.1 30.54±0.1 30.70±0.0
Law Random 32.60±0.3 33.52±0.5 34.28±0.4 34.80±0.5 35.21±0.5
NSP 32.89±0.1 34.18±0.0 34.73±0.1 35.16±0.3 35.53±0.0
IDDS 32.91±0.0 34.12±0.1 34.61±0.1 35.07±0.0 35.33±0.0
HUDS 33.76±0.0 34.39±0.2 34.89±0.1 35.17±0.0 35.62±0.1
Table 2: SacreBLEU scores with different active learning strategies for NMT across different domain datasets. BART-base model is used in all experiments. Active learning is done over ten iterations with 1000 sentences queried in each iteration. For each strategy, three independent runs are done with different seeds, and the mean SacreBLEU is reported, with the subscript representing the standard error. Best results are given in bold. HUDS consistently shows better performance compared to other AL strategies.